OpenStudy (anonymous):

Find the equation of the ellipse with the following properties: The ellipse with foci at (0, 6) and (0, -6); y-intercepts (0, 8) and (0, -8).

OpenStudy (anonymous):

u got y intercepts so u got 2b nd u got focii....so u got 2ae.... ok?

OpenStudy (anonymous):

distance b/n focus is 2ae...........nd distance b/n y intercepts is 2b nd b2=a2(1-e2) find b and a equation of ellipse is x2/a2 + y2/b2 = 1

OpenStudy (anonymous):

I still don`t get it..

OpenStudy (anonymous):

2b=(6 +6) b=6 nd 2ae =8 + 8 ae = 8